OpenStudy (anonymous):

a girl is y years old she is twice as old as her brother.Their combined ages total 16.Give an expression that shows their ages

OpenStudy (whpalmer4):

If the girl is \(y\) years old, and is twice as old as her brother, how would you write an expression for her brother's age?

OpenStudy (anonymous):

what will be the answer

OpenStudy (anonymous):

let brother age be x y=2x 2x-y=0 and x+y=16

OpenStudy (anonymous):

y = 2 b, b = y/2